A force of 25 N is applied along the handle of a gurney which is pushed along the hospital corridor.Find the horizontal and vertical component of this force when the handle is held 30 degres 40 degres and 60 degress
1
Expert's answer
2020-10-05T10:56:09-0400

1)


"F_x=F\cos{30}=25\cos{30}=21.65\ N\\F_y=F\sin{30}=25\sin{30}=12.5\ N"

2)


"F_x=F\cos{40}=25\cos{40}=19.15\ N\\F_y=F\sin{40}=25\sin{40}=16.07\ N"

3)


"F_x=F\cos{60}=25\cos{60}=12.5\ N\\F_y=F\sin{60}=25\sin{60}=21.65\ N"